# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 7. Exceptional set estimate: Upper bound I
Proposition 38.
Assume Conjecture 4 holds. For $10<a\leq k+$ 1 and $0<s$, and $A\subset\mathbb{F}_{p}^{k+1}$ with #$\#A\gtrsim p^{a}$, we have
(70) #$\#E_{s-\varepsilon}(A;k+1,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)},$
@@
Lemma 39.
Assume Conjecture 4 holds. Fix $1000<\varepsilon_{0}<1/$ 100. Let $00<a\leq k+1,s>$ 0 and $\sigma\geq 0$. Suppose the parameters satisfy $\mathbf{M}(a,s;k+1,k)+\varepsilon_{0}\leq k$ and $a-s+\varepsilon_{0}\leq\sigma\leq a$. Let $A\subset\mathbb{F}_{p}^{k+1}$ with #$\#A=p^{k+1}$. Define
@@
#$\#\mathcal{V}\lesssim_{\varepsilon_{0}}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
Remark 40.
The implicit constant in $\lesssim_{\varepsilon_{0}}$ can be made independent of $\sigma$, since we can apply the lemma to $\sigma\in\varepsilon_{0}^{2}\cdot\mathbb{N}$ which consists $O(\varepsilon_{0}^{-2})$ values.
@@
The proof of this part is by Furstenberg estimate. As in Definition 14, write
(71) $\begin{cases}a=m+\beta\\ s=l+\gamma.\end{cases}$
@@
Type 1 is trivial, since #$\#\mathcal{V}\leq p^{k}$ is always true. Type 4 is also easy, since in this case we have $\sigma\geq a-s+\varepsilon_{0}\geq 1+\varepsilon_{0}$, and hence $\mathcal{V}=\emptyset$. We only need to consider Type 2 and Type 3.
Suppose #$\#\mathcal{V}=p^{t}$, we want to show
@@
(72) $R(\beta,\gamma):=\begin{cases}\max\{2\gamma-(\beta+1),0\}&m=l+1,\gamma\in(\beta,1]\\ \max\{2\gamma-\beta,0\}&m=l,\gamma\in(0,\beta].\end{cases}$
@@
Note that $A$ contains a $(\sigma,t;k+1,1)$-Furstenberg set $\bigcup_{L\in\mathcal{V}}(L\cap A)$. Since Conjecture 4 holds, by Theorem 7, we have the lower bound:
(73) $a+\eta_{0}\geq\mathbf{F}(\sigma,t;k+1,1)=w+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\},$
@@
Since $s\leq l+1$ and $w\in\mathbb{N}$, we have $w\leq l$. Therefore, with the condition that $\mathbf{M}(a,s;k+1,k)\geq l$, we see that
@@
We just need to prove
@@
This is easy to verify by checking either in case $m=l+1,\gamma\in(\beta,1]$ or in case $m=l,\gamma\in(0,\beta]$. ∎
Proof of Proposition 38.
We will use the lemma that we just proved. We show that, if $A\subset\mathbb{F}_{p}^{k+1}$ with #$\#A\gtrsim p^{a}$, $E\subset G(1,\mathbb{F}_{p}^{k+1})$ with #$\#E>C_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)}$, and $p$ is big enough, then there exists $V\in E$ such that
#$\#\pi_{V}^{*}(A)>p^{s-\varepsilon}.$
Since #$\#E\leq p^{k}$, we have
@@
For each $V\in E$ and dyadic number $1\leq\mu\leq p$, define
@@
We see that $\{\mathbb{L}_{V,\mu}\}_{\mu}$ form a partition of the of lines that are parallel to $V$.
We choose $\varepsilon_{0}\ll\varepsilon$, for example, $\varepsilon_{0}=\varepsilon^{2}$. We will estimate the cardinality of the set $\bigcup_{V\in E}\mathbb{L}_{V,\mu}$ for $\mu\geq p^{a-s+\varepsilon_{0}}$. We write $\mu=p^{\sigma}$ with $a\geq\sigma\geq a-s+\varepsilon_{0}$. By Lemma 39, we have
#$\#\bigcup_{V\in E}\mathbb{L}_{V,\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
@@
#$\#E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+2\varepsilon_{0}}.$
@@
#$\#\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+3\varepsilon_{0}}.$
Therefore, #$\#E>\#\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$ when $p$ is large enough. We can find $V\in E\setminus\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$. In other words, there exists $V\in E$, such that for any $\mu=p^{\sigma}\geq p^{a-s+\varepsilon_{0}}$,
#$\#\mathbb{L}_{V,\mu}\leq p^{a-\sigma-\varepsilon_{0}}.$
@@
#$\#A=\sum_{\mu<p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L)+\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L).$
@@
This implies #$\#\pi_{V}^{*}(A)\gtrsim p^{s-\varepsilon_{0}}>p^{s-\varepsilon}$.
8. Exceptional set estimate: Upper bound II
@@
(74) $\sup_{A\subset\mathbb{F}_{p}^{n^{\prime}},\#A\sim p^{a}}\#E_{s-\varepsilon}(A;n^{\prime},k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n^{\prime},k)},$
@@
Lemma 42.
Assume (74) holds. Suppose $n,k\in\mathbb{N}$, $n\geq k+2$, $a\in(0,n]$, $s>0$ and $s\in(a-(n-k),\min\{a,k\}]$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with #$\#A\sim p^{a}$, there exist numbers $a_{1},s_{1}$, such that
(75) $\begin{cases}p^{\max\{0,a-1\}}\leq p^{a_{1}}\leq p^{\min\{n-1,a\}};\\ 1\leq p^{s_{1}}\leq p^{s};\end{cases}$
@@
(76) #$\#E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)}.$
@@
By properly choosing the coordinate, we can assume $\gtrsim 1$ fraction of the $(n-k)$-planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ transversely (at a $(n-k-1)$-plane). Since we are going to find upper bound for #$\#\mathcal{V}$, we can just assume all the planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ at a $(n-k-1)$-plane.
For each $W\in G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$, we define $\mathcal{V}_{W}:=\{V\in\mathcal{V}:V\cap(\mathbb{F}_{p}^{n-1}\times\mathbf{0})=W\}$. By pigeonholing, there exits $\mathcal{W}\subset G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with #$\#\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, #$\#\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $p^{t_{1}},p^{t_{2}}\geq 1$ are dyadic numbers so that
@@
By pigeonholing on #$\#A\cap(\mathbb{F}_{p}^{n-1}\times\xi)$ for $\xi\in\mathbb{F}_{p}$, we can find a dyadic number $1\leq p^{a_{1}}\leq p^{a}$ and $\Xi_{1}\subset\mathbb{F}_{p}$ so that #$\#A\cap(\mathbb{F}_{p}^{n-1}\times\xi)\sim p^{a_{1}}$ for any $\xi\in\Xi_{1}$ and
@@
Since #$\#\Xi_{1}\leq p$, we also have $p^{a_{1}}\gtrapprox p^{a-1}$.
@@
#$\#\pi_{W}^{*}(A_{\xi})\sim p^{s_{1}(\xi)},$
for #$\gtrapprox\#\mathcal{W}\sim p^{t_{1}}$ many $W\in\mathcal{W}$. Here, $s_{1}(\xi)\leq s-\varepsilon$ is because #$\#\pi_{W}^{*}(A_{\xi})\leq\#\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$.
@@
Define $B_{\xi}=\pi_{\mathbb{F}_{p}^{k+1}}(A_{\xi})$. We see that $B_{\xi}$ is a subset of $\mathbf{0}^{n-1-k}\times\mathbb{F}_{p}^{k+1}$ and #$\#B_{\xi}=\#\pi_{W}^{*}(A_{\xi})$.
@@
We can now do the estimate. By definition, #$\#\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$. Therefore,
@@
#$\#\pi_{\mathbb{F}_{p}^{k+1}}(A)\geq\#\pi_{\mathbb{F}_{p}^{k+1}}(\bigcup_{\xi\in\Xi}A_{\xi})=\#\pi^{*}_{W}(\bigcup_{\xi\in\Xi}A_{\xi})=\sum_{\xi\in\Xi}\#\pi_{W}^{*}(A_{\xi})\gtrapprox p^{a-a_{1}+s_{1}}.$
@@
#$\#\mathcal{V}\lessapprox p^{t_{1}+t_{2}}\lesssim_{\varepsilon}p^{\varepsilon/2+\mathbf{M}(a_{1},s_{1}+\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}-\varepsilon^{2},s-\varepsilon/2;k+1,k)}.$
@@
Recall $23\varepsilon^{2}<\varepsilon/$ 2 and note $\mathbf{M}(a,s;n,k)$ is decreasing in $a$. By Lemma 23, we have
@@
#$\#\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1}^{\prime}-\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}^{\prime},s;k+1,k)}.$
Next, we want to slightly increase $a_{1}$ to $a_{1}^{\prime}$ so that $p^{\min\{0,a-1\}}\leq p^{a_{1}^{\prime}}$. By the condition $p^{\min\{0,a-1\}}\lessapprox p^{a_{1}}$, we have $p^{\min\{0,a-1\}}\leq p^{a_{1}+\varepsilon^{2}}$ when $p$ is large enough. We just need to let $a_{1}^{\prime}=\min\{a_{1}+\varepsilon^{2},\min\{0,a-1\}\}$. Noting $a_{1}^{\prime}\leq a_{1}+\varepsilon^{2}$ and by Lemma 23 again, we have
@@
#$\#\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(a-a_{1}+s_{1},s;k+1,k)}.$
Lemma 43.
@@
(77) $\begin{cases}\max\{0,a-1\}\leq a_{1}\leq\min\{n-1,a\};\\ 0\leq s_{1}\leq s.\end{cases}$
@@
(78) $\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)\leq\mathbf{M}(a,s;n,k).$
@@
Just need to note $m\geq l$ and $2\gamma_{1}\leq 2$.
@@
Next we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. We claim we have
(79) $\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$
We just consider when $\gamma_{1}+\beta-\beta_{1}\geq\gamma$. We note $\gamma_{1}+\beta-\beta_{1}\in(0,2)$. If $\gamma_{1}+\beta-\beta_{1}\in(0,1]$, then we use Type 3 in Definition 14, which gives the desired formula. If $\gamma_{1}+\beta-\beta_{1}\in(1,2]$, we use Type 2, which gives $\mathbf{M}(l+1+(\gamma_{1}+\beta-\beta_{1}-1),l+\gamma;k+1,k)=k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}$, which is the same.
@@
In this case, by Lemma 24,
@@
(We use $\leq$ instead of $=$, since when $m=l$, $\mathbf{M}(a_{1},s_{1};n-1,k)=k(n-1-k)$.)
@@
Case 3: $m_{1}=m-1,l_{1}\leq l-2$ The proof is similar to Case 1. By Lemma 24,
@@
Just need to note $m-l\geq 1$ and $2\gamma_{1}\leq 2$.
@@
Next we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. As is Case 2.1,
(80) $\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$
@@
When $\gamma_{1}+\beta-\beta_{1}<\gamma$, we just need to prove
@@
When $\gamma_{1}+\beta-\beta_{1}\geq\gamma$, we just need to prove
@@
Just need to check three scenarios:
@@
Also, we can assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it =absent=$\infty$= - ∞.
@@
When $m=l$, we just need to prove
@@
We can also assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it =absent=$\infty$= - ∞.
@@
(81) $\gamma>\gamma_{1}+\beta-\beta_{1},$
otherwise this is $\infty$- ∞.
@@
(82) $\gamma>\gamma_{1}+\beta-\beta_{1},$
otherwise this is .$\infty$.- ∞ .

### v2

@@ 7 Exceptional set estimate: Upper bound I
Proposition 7.1.
Assume Conjecture 1.5 holds. For $0<a\leq k+1$ and $0<s$, and $A\subset\mathbb{F}_{p}^{k+1}$ with $A\gtrsim p^{a}$, we have
$E_{s-\varepsilon}(A;k+1,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)},$ (70)
@@
Lemma 7.2.
Assume Conjecture 1.5 holds. Fix $0<\varepsilon_{0}<1/100$. Let $0<a\leq k+1,s>0$ and $\sigma\geq 0$. Suppose that the parameters satisfy $\mathbf{M}(a,s;k+1,k)+\varepsilon_{0}\leq k$ and $a-s+\varepsilon_{0}\leq\sigma\leq a$. Let $A\subset\mathbb{F}_{p}^{k+1}$ with $A=p^{k+1}$. Define
@@
$\mathcal{V}\lesssim_{\varepsilon_{0}}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
Remark 7.3.
The implicit constant in $\lesssim_{\varepsilon_{0}}$ can be made independent of $\sigma$, since we can apply the lemma to $\sigma\in\varepsilon_{0}^{2}\cdot\mathbb{N}$ which consists of $O(\varepsilon_{0}^{-2})$ values.
@@
The proof of this part is by a Furstenberg estimate. As in Definition 2.4, write
$\begin{cases}a=m+\beta\\ s=l+\gamma.\end{cases}$ (71)
@@
Type 1 is trivial, since $\mathcal{V}\leq p^{k}$ is always true. Type 4 is also easy, since in this case we have $\sigma\geq a-s+\varepsilon_{0}\geq 1+\varepsilon_{0}$, and hence $\mathcal{V}=\emptyset$. We only need to consider Type 2 and Type 3.
Suppose $\mathcal{V}=p^{t}$, we want to show
@@
$R(\beta,\gamma):=\begin{cases}\max\{2\gamma-(\beta+1),0\}&m=l+1,\gamma\in(\beta,1]\\ \max\{2\gamma-\beta,0\}&m=l,\gamma\in(0,\beta].\end{cases}$ (72)
@@
Note that $A$ contains a $(\sigma,t;k+1,1)$-Furstenberg set $\bigcup_{L\in\mathcal{V}}(L\cap A)$. Since Conjecture 1.5 holds, by Theorem 1.8, we have the lower bound:
$a+\eta_{0}\geq\mathbf{F}(\sigma,t;k+1,1)=w+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\},$ (73)
@@
Since $s\leq l+1$ and $w\in\mathbb{N}$, we have $w\leq l$. Therefore, with the condition $\mathbf{M}(a,s;k+1,k)\geq l$, we see that
@@
We only need to prove
@@
This is easy to verify by checking either in the case $m=l+1,\gamma\in(\beta,1]$ or in the case $m=l,\gamma\in(0,\beta]$. ∎
Proof of Proposition 7.1.
We will use the lemma that we just proved. We show that if $A\subset\mathbb{F}_{p}^{k+1}$ with $A\gtrsim p^{a}$, $E\subset G(1,\mathbb{F}_{p}^{k+1})$ with $E>C_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)}$, and $p$ is big enough, then there exists $V\in E$ such that
$\pi_{V}^{*}(A)>p^{s-\varepsilon}.$
Since $E\leq p^{k}$, we have
@@
For each $V\in E$ and a dyadic number $1\leq\mu\leq p$, define
@@
We see that $\{\mathbb{L}_{V,\mu}\}_{\mu}$ form a partition of the lines parallel to $V$.
We choose $\varepsilon_{0}\ll\varepsilon$, for example, $\varepsilon_{0}=\varepsilon^{2}$. We will estimate the cardinality of the set $\bigcup_{V\in E}\mathbb{L}_{V,\mu}$ for $\mu\geq p^{a-s+\varepsilon_{0}}$. We write $\mu=p^{\sigma}$ with $a\geq\sigma\geq a-s+\varepsilon_{0}$. By Lemma 7.2, we have
$\bigcup_{V\in E}\mathbb{L}_{V,\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
@@
$E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+2\varepsilon_{0}}.$
@@
$\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+3\varepsilon_{0}}.$
Therefore, $E>\#\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$ when $p$ is large enough. We can find $V\in E\setminus\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$. In other words, there exists $V\in E$, such that for any $\mu=p^{\sigma}\geq p^{a-s+\varepsilon_{0}}$,
$\mathbb{L}_{V,\mu}\leq p^{a-\sigma-\varepsilon_{0}}.$
@@
$A=\sum_{\mu<p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L)+\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L).$
@@
This implies $\pi_{V}^{*}(A)\gtrsim p^{s-\varepsilon_{0}}>p^{s-\varepsilon}$.
8 Exceptional set estimate: Upper bound II
@@
$\sup_{A\subset\mathbb{F}_{p}^{n^{\prime}},\#A\sim p^{a}}\#E_{s-\varepsilon}(A;n^{\prime},k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n^{\prime},k)},$ (74)
@@
Lemma 8.2.
Assume (74) holds. Suppose $n,k\in\mathbb{N}$, $n\geq k+2$, $a\in(0,n]$, $s>0$ and $s\in(a-(n-k),\min\{a,k\}]$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\sim p^{a}$, there exist numbers $a_{1},s_{1}$, such that
$\begin{cases}p^{\max\{0,a-1\}}\leq p^{a_{1}}\leq p^{\min\{n-1,a\}};\\ 1\leq p^{s_{1}}\leq p^{s};\end{cases}$ (75)
@@
$E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)}.$ (76)
@@
By properly choosing the coordinate, we can assume $\gtrsim 1$ fraction of the (nk)(n-k)-planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ transversely (at a (nk1)(n-k-1)-plane). Since we are going to find upper bound for $\mathcal{V}$, we can just assume all the planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ at a (nk1)(n-k-1)-plane.
For each $W\in G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$, we define $\mathcal{V}_{W}:=\{V\in\mathcal{V}:V\cap(\mathbb{F}_{p}^{n-1}\times\mathbf{0})=W\}$. By pigeonholing, there exits $\mathcal{W}\subset G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $p^{t_{1}},p^{t_{2}}\geq 1$ are dyadic numbers so that
@@
By pigeonholing on $A\cap(\mathbb{F}_{p}^{n-1}\times\xi)$ for $\xi\in\mathbb{F}_{p}$, we can find a dyadic number $1\leq p^{a_{1}}\leq p^{a}$ and $\Xi_{1}\subset\mathbb{F}_{p}$ so that $A\cap(\mathbb{F}_{p}^{n-1}\times\xi)\sim p^{a_{1}}$ for any $\xi\in\Xi_{1}$ and
@@
Since $\Xi_{1}\leq p$, we also have $p^{a_{1}}\gtrapprox p^{a-1}$.
@@
$\pi_{W}^{*}(A_{\xi})\sim p^{s_{1}(\xi)},$
for $\gtrapprox\#\mathcal{W}\sim p^{t_{1}}$ many $W\in\mathcal{W}$. Here, $s_{1}(\xi)\leq s-\varepsilon$ is because $\pi_{W}^{*}(A_{\xi})\leq\#\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$.
@@
Define $B_{\xi}=\pi_{\mathbb{F}_{p}^{k+1}}(A_{\xi})$. We see that $B_{\xi}$ is a subset of $\mathbf{0}^{n-1-k}\times\mathbb{F}_{p}^{k+1}$ and $B_{\xi}=\#\pi_{W}^{*}(A_{\xi})$.
@@
We can now do the estimate. By definition, $\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$. Therefore,
@@
$\pi_{\mathbb{F}_{p}^{k+1}}(A)\geq\#\pi_{\mathbb{F}_{p}^{k+1}}(\bigcup_{\xi\in\Xi}A_{\xi})=\#\pi^{*}_{W}(\bigcup_{\xi\in\Xi}A_{\xi})=\sum_{\xi\in\Xi}\#\pi_{W}^{*}(A_{\xi})\gtrapprox p^{a-a_{1}+s_{1}}.$
@@
$\mathcal{V}\lessapprox p^{t_{1}+t_{2}}\lesssim_{\varepsilon}p^{\varepsilon/2+\mathbf{M}(a_{1},s_{1}+\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}-\varepsilon^{2},s-\varepsilon/2;k+1,k)}.$
@@
Recall $3\varepsilon^{2}<\varepsilon/2$ and note $\mathbf{M}(a,s;n,k)$ is decreasing in $a$. By Lemma 2.15, we have
@@
$\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1}^{\prime}-\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}^{\prime},s;k+1,k)}.$
Next, we want to slightly increase $a_{1}$ to $a_{1}^{\prime}$ so that $p^{\min\{0,a-1\}}\leq p^{a_{1}^{\prime}}$. By the condition $p^{\min\{0,a-1\}}\lessapprox p^{a_{1}}$, we have $p^{\min\{0,a-1\}}\leq p^{a_{1}+\varepsilon^{2}}$ when $p$ is large enough. We just need to let $a_{1}^{\prime}=\min\{a_{1}+\varepsilon^{2},\min\{0,a-1\}\}$. Noting $a_{1}^{\prime}\leq a_{1}+\varepsilon^{2}$ and by Lemma 2.15 again, we have
@@
$\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(a-a_{1}+s_{1},s;k+1,k)}.$
Lemma 8.3.
@@
$\begin{cases}\max\{0,a-1\}\leq a_{1}\leq\min\{n-1,a\};\\ 0\leq s_{1}\leq s.\end{cases}$ (77)
@@
$\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)\leq\mathbf{M}(a,s;n,k).$ (78)
@@
We just need to note $m\geq l$ and $2\gamma_{1}\leq 2$.
@@
Next, we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. We claim that we have
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$ (79)
We just consider when $\gamma_{1}+\beta-\beta_{1}\geq\gamma$. We note $\gamma_{1}+\beta-\beta_{1}\in(0,2)$. If $\gamma_{1}+\beta-\beta_{1}\in(0,1]$, then we use Type 3 in Definition 2.4, which gives the desired formula. If $\gamma_{1}+\beta-\beta_{1}\in(1,2]$, we use Type 2, which gives $\mathbf{M}(l+1+(\gamma_{1}+\beta-\beta_{1}-1),l+\gamma;k+1,k)=k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}$, which is the same.
@@
In this case, by Lemma 2.16,
@@
(We use $\leq$ instead of ==, since when $m=l$, $\mathbf{M}(a_{1},s_{1};n-1,k)=k(n-1-k)$.)
@@
Case 3: $m_{1}=m-1,l_{1}\leq l-2$The proof is similar to Case 1. By Lemma 2.16,
@@
We just need to note $m-l\geq 1$ and $2\gamma_{1}\leq 2$.
@@
Next, we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. As is Case 2.1,
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$ (80)
@@
When $\gamma_{1}+\beta-\beta_{1}<\gamma$, we only need to prove
@@
When $\gamma_{1}+\beta-\beta_{1}\geq\gamma$, we only need to prove
@@
We just need to check three scenarios:
@@
Also, we can assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it $\infty$.
@@
When $m=l$, we only need to prove
@@
We can also assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it $\infty$.
@@
$\gamma>\gamma_{1}+\beta-\beta_{1}.$ (81)
Otherwise, this is -$\infty$.
@@
$\gamma>\gamma_{1}+\beta-\beta_{1}.$ (82)
Otherwise, this is $\infty.$